A quadratic function has zeros x=2/3 and x=-5. What two possible equations can solve this function?
Answer by stanbon(75887)

A quadratic function has zeros x=2/3 and x=-5. What two possible equations can solve this function?
-----
Since x = 2/3 is a zero, 3x-2 is a factor
Since x = -5 is a zero, x+5 is a factor
---
Equation Form y = a(3x-2)(x+5) where a is a non-zero Real Number.
----
Answer:: y = 2(3x-2)(x+5) or y = 3(3x-2)(x+5)
---------
Cheers,
Stan H.
--------